JiscMail Logo
Email discussion lists for the UK Education and Research communities

Help for UKHEPGRID Archives


UKHEPGRID Archives

UKHEPGRID Archives


UKHEPGRID@JISCMAIL.AC.UK


View:

Message:

[

First

|

Previous

|

Next

|

Last

]

By Topic:

[

First

|

Previous

|

Next

|

Last

]

By Author:

[

First

|

Previous

|

Next

|

Last

]

Font:

Proportional Font

LISTSERV Archives

LISTSERV Archives

UKHEPGRID Home

UKHEPGRID Home

UKHEPGRID  February 2008

UKHEPGRID February 2008

Options

Subscribe or Unsubscribe

Subscribe or Unsubscribe

Log In

Log In

Get Password

Get Password

Subject:

Minutes of the 291st GridPP PMB meeting

From:

Tony Doyle <[log in to unmask]>

Reply-To:

Tony Doyle <[log in to unmask]>

Date:

Fri, 15 Feb 2008 10:43:40 +0000

Content-Type:

MULTIPART/MIXED

Parts/Attachments:

Parts/Attachments

TEXT/PLAIN (20 lines) , 080211.txt (1 lines)

Dear All,

     Please find attached the latest GridPP Project Management Board 
Meeting minutes. The latest minutes can be found each week in:

http://www.gridpp.ac.uk/php/pmb/minutes.php?latest

as well as being listed with other minutes at:

http://www.gridpp.ac.uk/php/pmb/minutes.php

Cheers, Tony
________________________________________________________________________
Prof. A T Doyle, FInstP FRSE                       GridPP Project Leader
Rm 478, Kelvin Building                      Telephone: +44-141-330 5899
Dept of Physics and Astronomy                  Telefax: +44-141-330 5881
University of Glasgow                   EMail: [log in to unmask]
G12 8QQ, UK                 Web: http://ppewww.physics.gla.ac.uk/~doyle/
________________________________________________________________________


GridPP PMB Minutes 291 - 11th February 2008 =========================================== Present: Tony Doyle, Sarah Pearce, Roger Jones, Stephen Burke, David Britton, Steve Lloyd, John Gordon, Jeremy Coles, Peter Clarke, Glenn Patrick, Andrew Sansum, Dave Colling, Suzanne Scott (Minutes) Apologies: David Kelsey, Tony Cass, Robin Middleton, Neil Geddes SI-1 Dissemination Officer's Report ------------------------------------ SP reported two news items, one at draft stage regarding UKQCD on the PIPSS award relating to biotechnology, and the other relating to Neasan O'Neill currently attending the EGEE User Forum in France with the UKI stand. SP reported that Neasan O'Neill had met with A McNab regarding the GridPP Website initial beta release at the end of the month, discussing the new design and the background changes. There would be a test opportunity for the Website in March. If anyone wished anything included could they let SP know. Neasan O'Neill was also currently gathering updates for posters in relation to IoP HEPP Group meeting - these were generally updated once per year. It was noted that some were being omitted. AS and JC provided an update on the recent outage at the Tier-1 and its effect on Grid Production. The discussion would be continued at next week's PMB as one input to overall disaster planning. SI-2 Tier-1 Manager's Report ----------------------------- AS reported as follows: 3) While working with one resiliant transformer down the other transformer failed. We have not yet received a post-mortem analysis from building services - although the failure was probably related to the ongoing work. Power failed at about 12:10 on 7th February and was restored about an hour later. Most critical national and global core services were re-established later that day. The GOCDB was not available until the Friday morning after problems rolling the database forward from the Oracle journal. The CASTOR and Grid batch services were available late on Friday although full batch capacity did not appear until Saturday. Notification was problematic with no power/phones in the building and no means if issuing a broadcast (GOCDB down). Notification to the VOs was achieved within 30-60 minutes (ditto the PMB) and WLCG was informed by 14:30. 4) We have seen a couple of incidents where drive failures appear to have been handled incorrectly by the RAID controller. It appears likely that this is related to the backplane issues identified below. This item is closed unless we experience a recurrence. 5) On 31st January evening a disk server failed and smoke/fumes from the failure trigger the fire alarm. Investigation showed that a hole had been burnt through the disk backplane. This is the second failure to trigger the fire alarm and is being treated as a safety matter. We have received good support from the supplier and manufacturer and have now received a written report which indicates a batch-related problem with the server backplane. Although the supplier's assesment is that there is no significant risk to safety, our safety review concluded that we must address this problem as soon as practical. The supplier should have replacement backplanes early this week (they should have arrived in the UK on Friday) and they are reviewing what staff effort they have to carry out an urgent intervention. It probably will take 20-30 minutes per server and we have 86 servers to repair. It is likely that we will need to announce a lengthy downtime for mid/later this week but have yet to decide if we will take the whole service down at once or intervene VO by VO. AS reported that an analysis by the supplier had shown that the outage was caused by a printed circuit board and alignment problems with layers and crossover - this caused resistors to overheat. There had been three board failures to date since last August '07, but not all of them had set-off the fire alarms. AS was unsure as to exactly how much downtime was required. TD advised that doing it by VO seemed a sensible approach. AS confirmed that once he had more definitive information from the supplier as to what they can do, and staff availability to carry out the work etc, he would iterate with the experiments and would make a judgement about downtime at that point. AS further reported that as of today, all was working except the GOCDB, and there had also been a further CASTOR issue over the weekend. Later this week a retrospective assessment would be carried out. AS noted that volume of equipment was an issue, also the streamlining of staff in such an instance. Notification had also been an issue - AS had managed to contact TD and also Graeme Stewart regarding dissemination of the outage. AS had been advised to contact the CERN line 5501 should anything similar happen again. AS confirmed he would document all of the issues. TD asked JC to comment on the communication channels. JC noted that he had not yet done a retrospective analysis but would do that this week, however he had received five notifications within the hour, and will examine the impact of the outage. AS and JC would iterate on Thursday afternoon with a view to reporting back to the PMB next Monday. 1) Tenders: a) Disk tender - supplier load test completed. Our 28 day load test has not started and is now running late. We plan to start it this week, but have experienced disruption due to power and hardware problems. b) CPU tender - order placed and scheduled for delivery 28 February. Suppliers may deliver 1-2 weeks early. It will probably not be possible to complete the full 28-day acceptance test before it is necessary to pay the bill in this financial year. Once we have 1-2 weeks load test results the PMB will be asked to approve payment. c) Tape drive purchase - six tape drives have been received (5 drives are currently being borrowed). Tape servers must be ordered shortly. d) Non-Capacity hardware order has been placed. Delivery is expected to be 1-2 weeks later than the CPU delivery. e) Oracle server hardware upgrade order has been placed. f) An order for a 32 port non-blocking 10Gb switch has been placed. Delivery is expected in mid-March. 2) Work on the power supply is nearly complete. Switchover back to two supplies is scheduled for 13:00 today (Monday). Service: 1) SAM availability for last week was 75% - main problem was the power failure, although site DNS problems also impacted the service on Wednesday night. 2) CASTOR: a) ATLAS CMS and LHCB were fully configured for CCRC08. Alice have a disk allocation and have started to attend the CASTOR teleconference. Work on an Alice disk pool has now started. b) One of the CASTOR endpoints (ralsrmd) has been down since Saturday. A restart was attempted on Saturday but failed and the problem is still under investigation. 3) SL4 Migration - The SL4 UI build is now working and a test system was available. A minor problem was found and a new build was scheduled for last Thursday, unfortunately this did not happen and will now be completed this week. 4) At the PMB face to face an issue was raised about number of hosts with 2GB/core available. There are about 400 job slots guaranteed to have at least 2GB/core at their disposal (about half the farm by KSI2K). The remaining servers have 1GB/core and may be able to run 1.5GB jobs if there is a mix of jobs with lower memory requirements. DC noted that requirements had been growing over the past few months - very little was use of Tier-1, but was mostly scheduled reconstruction. AS noted that the older kit was about half of farm capacity, and was due to be phased-out. TD asked whether AS was now formally in 'recovered disaster' mode? AS said that this classification was unclear at present. TD advised that if there were to be imminent turn-off of disks this week - users needed to know. JG advised that the incident could be classed as not a disaster, but a major incident. TD suggested that users do need to know about significant downtime expected - and this would need to be disseminated. AS confirmed that once he had definitive feedback from the supplier about manpower and timescales, he would advertise any potential downtime. 5) VO boxes - We are working with Alice on an upgrade to SL4 - We have arranged with CMS to split the VO box into two, one running Frontier - the other Phedex. This will be carried out with a planned upgrade to Phedex. 6) AFS There have been performance/availability problems with AFS which have been traced to a misconfigured firewall at a remote site. This has been severely impacting Babar but has now been resolved. Progress to Grid Only Access ============================ This standing item documents the status of work towards achieving GRIDPP milestone 0.18 "Access to Tier-1 resources by Grid Interfaces Only" 1) We have a list of users allowed to submit via qsub. When non-Grid submission is re-instated only this list will be used. SI-3 Production Manager's Report --------------------------------- JC reported as follows: 1) The power cut that hit RAL last week took out the GOCDB. This has led to a number of observations including that the Broadcast system in some way depends on the GOCDB being available. We will review how communications then progressed to develop our disaster planning. The GOCDB problem was traced to the NGS Oracle Database which supports the GOCDB being down. This has now been fixed. 2) UK CA certificates were taken out of CERN VOMS last week as it was thought that they had been revoked. A user ticket hinted that a specific problem being experienced might be due to old CA DN information still being present and certificates based on it were then removed without further investigation - i.e. the old root certificate was suspected compromised. This resulted in the majority of UK issued certificates not working for a period of several hours, and until all users had been registered with both issuer names. JG noted that CERN hadn't kept their VOMS certificates up-to-date. TD asked what the Policy was - there had to be a chain of trust from the DN to the CA. JG would raise this as an item at the MB. JC confirmed he would put it on the Ops meeting Agenda. 3) CCRC08 February run has started. There was a meeting at CERN last Tuesday to discuss the direction. ATLAS intends to use T2s from this week. One question they face is how to enforce analysis at T2s only. The questions on what space tokens to use for each of the experiments now seems clearer. CCRC progress is being logged in e-logbooks here: https://prod-grid-logger.cern.ch/elog/CCRC'08+Observations/. Other links of interest will be found here: http://tinyurl.com/2tk5q9. TD reported that a small group now existed, comprising TD, RJ, Alan Barr and Dan Tovey, which had been working on the GridPP wiki regarding ATLAS FDR plans, showing a breakdown of sites and available streams. TD advised that the course of action might be to limit who can submit jobs to the Tier-1 or delete the AOD at the Tier-1, however a definitive solution was not yet apparent - it was noted that FDR tests the computing model, not just the resources. See https://www.gridpp.ac.uk/wiki/AtlasFdr1 4) The topic of SL4 was discussed at the GDB last week. The majority of UK sites now have plans to move and some have started asking about SL5. JG noted that SL5 had been discussed at the MB - the PMB should consult TC's slides as well as the conclusions outlined in the Minutes. TD sent round a url in the chat window relating to the SL4 discussion. https://twiki.cern.ch/twiki/pub/LCG/MbMeetingsMinutes/LCG_Management_Board_2008_02_05.htm "Post SLC4 options at CERN - Tony Cass" 5) GridIreland colleagues have now created their own blog (http://gridirelandops.blogspot.com/) and it is being aggregated along with our GridPP blogs. 6) The overall job load is up in the last week. Today CPU is running at roughly 55%. The UK SAM test average is down at 81% for the last week. Meetings A) There is an EGEE User Forum in Clermont-Ferrand this week (11th-14th): http://www.eu-egee.org/egee_events/userforum/3-user-forum/ SI-4 LCG Management Board Report --------------------------------- TD reported that an issue under discussion at present was user accounting and its impact within the experiments. There was an agreement that an interim policy, proposed by JG, needs to be implemented and that everyone needs to know how data is accounted for. JG noted that a prototype portal had been published to show what data exists. TD reported that there was a summary available of the discussion on CCRC; there had also been a discussion of the CASTOR metrics used to clarify how well the system was working internally; SL4 had been discussed along with what should happen afterwards and the timescale proposed. Other discussions covered VO box support - this issue would be addressed this week - there had been a presentation from Alice and also a review of the applications area. See https://twiki.cern.ch/twiki/pub/LCG/MbMeetingsMinutes/LCG_Management_Board_2008_02_05.htm SI-5 Documentation Officer's Report ------------------------------------ SB noted that items previously discussed were all ongoing. AOCB ==== GP reported that the LHCb software course had been delayed and might not happen until the next financial year. He asked whether it would be possible to carry forward the 6k allocated by GridPP for this. TD noted that information was still awaited from RM regarding action 290.20 relating to travel figures and information. DB advised that if the course was delayed then in principle it would still be supported, but he noted that if there were to be financial difficulties then and that expenditure would be considered along with all other expenditure. TD reported that the announcements from STFC were expected at the end of February, following which there would be a three-week consultation in March. REVIEW OF ACTIONS ================= 272.4 AS to check the current Tier-1 disaster recovery plan and circulate the existing version to the PMB. It was reported that this document does not exist, but it was planned to have one in the longer term. TD would incorporate in v0.4 anything that AS considered relevant. AS will check and advise additions. Done, item closed. 277.2 DN to provide an update and re-evaluation of CMS/CASTOR deliverables. TD advised that there was a CMS/CASTOR document on deliverables which should be revised in light of the December '07 tests. DC to take the token for this now and iterate with DN. DC reported that the document would be sent out this week. 277.5 Disaster Recovery 'Team B': SB, JC, TD, SP, DB to analyse the wider issues of disaster planning, mapped to the experiments' lists, and this work would include Project Management. A Recovery Plan was required. It was agreed that JC was in charge of this and the experiment input relating to subsets of the disaster plan. SB/JC to progress. This was now being dealt with via F2F actions. Done, item closed. 277.8 User Experience 'Team C': SB, SP, SL, with input from JC to deal with the issue of user experience and design of an easily-found lookup facility for grid error messages. SL reported that he had started the ATLAS wiki page and would circulate the url. SB was leading this with inputs from SP, SL and JC where needed. A new simple summary was required of all areas available plus a lookup/links facility, for the OC to review. This would include a list of most recent types of problems (possibly a 'top 12' for users - what the error means and the course of action to follow). SB to progress this. Ongoing. 280.7 JC to mention the issues (when approached by a VO with regard to joining) of the 'standard' 6-month introduction period, following which the VO must set-up something specific to them, if appropriate. This was discussed at DTeam. JC to email GridPP VO members if possible - ongoing. This was a standing action - JC had discussed it with the Tier-2 Co-ordinators in relation to VO members. JC to send email. Ongoing. 280.8 JG to investigate the UKI ROC website - any change/progress, and report-back. SB to iterate with JG in order to sign-off this item next week. Done, item closed. 282.2 SP to progress the Project Map using the T1 service areas and input from the meeting. Done, item closed. 282.6 JC and SB to progress existing 'disaster planning' template for next F2F meeting on 1st Feb. Involve experiments as necessary. This was a follow-up from the last F2F, and was to be distinguished from 277.5 action which is a longer-term one relating to the OC. Done, item closed. 289.1 AS to provide an analysis of the ATLAS disk server failures on the RAID controller. It was noted that this may be due to the backplanes problem, and was considered non-catastrophic - outcome awaited of the investigation. Done, item closed. 289.2 DC to check current situation regarding gLite WMS and SL4 - current status to be conveyed to DTeam. Ongoing. 289.3 JC to check the VOMS/-skipcacheck issue (in relation to UK CA certificate change) with Jens Jensen and raise the issue at an Operations meeting. This was raised via a ticket and will be discussed today at the Ops meeting. Done, action closed. 289.4 SP to speak to the KT person at STFC who assisted with the PIPSS case, to help with the post-competitive phase (in relation to EGEE only providing support to pre-competitive startup). SP to involve NG. SP had emailed and was awaiting a reply. Done, item closed. 290.1 JC to write down membership of DTeam. 290.2 RJ, DC and GP to nominate experiment user representatives for the Deployment Board. 290.3 SL and DB to review the Tier-1 Board Terms of Reference and see what could be formally incorporated into the new Deployment Board Terms of Reference. 290.4 AS and JG to iterate regarding what could replace the Tier-1 Board. 290.5 All: to check their individual roles as outlined and advise DB of any required changes. Ongoing. 290.6 TD to contact Iain and suggest that the GDB and MB roles change as at 1st April 2008. It was noted that DB would take over on the MB from 1st April. Done, item closed. 290.7 AS to provide numbers in the Quarterly Report for the Tier-1 as per the ones provided for Tier-2. 290.8 AS/SP to iterate regarding the financial summary in the Quarterly Reporting (eg: Outturn figures). 290.9 Quarterly Report for Tier-2 staff to be compiled by the Production Manager. 290.10 TD as Technical Director to provide a report showing effort figures; milestones & metrics; and a table of posts showing Technical Support. 290.11 DB to progress the situation at Manchester. 290.12 GP/SB/DC to define these Support posts and ensure they form a comprehensive basis for user support (both documentation and Grid access assistance), overseen by the UB Chair. 290.13 DB to complete the document re Reporting and Reporting Routes relating to staff, and circulate it, thereafter it would be posted on the website as a record. 290.14 RM to circulate the EGI Workshop Agenda. 290.15 JG to check with Malcolm Atkinson re attending the next EGI workshop in Rome (March). 290.16 NG noted that he had provided a draft paper relating to the end of EGEE III but would add information that addressed the period beyond 2011 and re-circulate. 290.17 Re the Project Map, SP would look at the EGI wiki, and NG would consider more inputs relating to box 6.2. 290.18 Regarding the LCG box on the Project Map, SP to iterate with TC and bring this issue back to the PMB. 290.19 DB/SP to progress the details of the Project Map over the next few months, cross-checking that all elements are incorporated, including strategic priorities and staffing. To be completed before the next Oversight Committee. It was agreed to move this to the 'inactive' category. 290.20 RM to provide more detailed figures on travel expenditure - broad-brush percentages would assist with decisions re travel in GridPP3. 290.21 SS to hand-out travel forms at Dublin ('overseas' claim on web to be submitted as 'actuals' and should be submitted before the end of March 2008). 290.22 AS to get back to RJ regarding job slots at the Tier-1. Done, item closed. 290.23 AS/JC to iterate on the Disaster Recovery template and remove capturable items that were considered to be minor. 290.24 JC to progress his suggested template to use when a crisis occurs - to be revisited subsequently at a PMB. ACTIONS AS AT 11.02.08 ====================== 277.2 DN to provide an update and re-evaluation of CMS/CASTOR deliverables. TD advised that there was a CMS/CASTOR document on deliverables which should be revised in light of the December '07 tests. DC to take the token for this now and iterate with DN. DC reported that the document would be sent out this week. 277.8 User Experience 'Team C': SB, SP, SL, with input from JC to deal with the issue of user experience and design of an easily-found lookup facility for grid error messages. SL reported that he had started the ATLAS wiki page and would circulate the url. SB was leading this with inputs from SP, SL and JC where needed. A new simple summary was required of all areas available plus a lookup/links facility, for the OC to review. This would include a list of most recent types of problems (possibly a 'top 12' for users - what the error means and the course of action to follow). SB to progress this. 280.7 JC to mention the issues (when approached by a VO with regard to joining) of the 'standard' 6-month introduction period, following which the VO must set-up something specific to them, if appropriate. This was discussed at DTeam. JC to email GridPP VO members if possible - ongoing. This was a standing action - JC had discussed it with the Tier-2 Co-ordinators in relation to VO members. JC to send email. 289.2 DC to check current situation regarding gLite WMS and SL4 - current status to be conveyed to DTeam. 290.1 JC to write down membership of DTeam. 290.2 RJ, DC and GP to nominate experiment user representatives for the Deployment Board. 290.3 SL and DB to review the Tier-1 Board Terms of Reference and see what could be formally incorporated into the new Deployment Board Terms of Reference. 290.4 AS and JG to iterate regarding what could replace the Tier-1 Board. 290.5 All: to check their individual roles as outlined and advise DB of any required changes. DB advised that he required input by next Monday 18th. 290.7 AS to provide numbers in the Quarterly Report for the Tier-1 as per the ones provided for Tier-2. 290.8 AS/SP to iterate regarding the financial summary in the Quarterly Reporting (eg: Outturn figures). 290.9 Quarterly Report for Tier-2 staff to be compiled by the Production Manager. 290.10 TD as Technical Director to provide a report showing effort figures; milestones & metrics; and a table of posts showing Technical Support. 290.11 DB to progress the situation at Manchester. 290.12 GP/SB/DC to define the portal and documentation Support posts and ensure they form a comprehensive basis for user support (both documentation and Grid access assistance), overseen by the UB Chair. 290.13 DB to complete the document re Reporting and Reporting Routes relating to staff, and circulate it, thereafter it would be posted on the website as a record. 290.14 RM to circulate the EGI Workshop Agenda. 290.15 JG to check with Malcolm Atkinson re attending the next EGI workshop in Rome (March). 290.16 NG noted that he had provided a draft paper relating to the end of EGEE III but would add information that addressed the period beyond 2011 and re-circulate. 290.17 Re the Project Map, SP would look at the EGI wiki, and NG would consider more inputs relating to box 6.2. 290.18 Regarding the LCG box on the Project Map, SP to iterate with TC and bring this issue back to the PMB. 290.20 RM to provide more detailed figures on travel expenditure - broad-brush percentages would assist with decisions re travel in GridPP3. 290.21 SS to hand-out travel forms at Dublin ('overseas' claim on web to be submitted as 'actuals' and should be submitted before the end of March 2008). 290.23 AS/JC to iterate on the Disaster Recovery template and remove capturable items that were considered to be minor. 290.24 JC to progress his suggested template to use when a crisis occurs - to be revisited subsequently at a PMB. 291.01 AS and JC to iterate on Thursday afternoon with a view to reporting back on the recent Tier-1 outage to the PMB next Monday. 291.02 JG to raise the issue of UK CA certificates being taken out of CERN VOMS, as an item at the MB. JC confirmed he would put it on the Ops meeting Agenda. INACTIVE CATEGORY ================= 271.1 PMB to examine the issue of fibre breakage and outages, CERN-RAL OPN link, in one year's time, when actual data on breakages is available. Due date would be September '08. 271.3 Re CERN-RAL OPN link breakage and backup generally, PC to oversee the issue and collate info so that the PMB have something to revisit in one year's time. Due date September '08. It was noted that PC would circulate a revised document after discussion with ATLAS (RJ/PC/DN to iterate). 282.8 RM to monitor how R-GMA and networking issues impact on GridPP as matters progress. RM advised that this item should be moved to the 'inactive' category as it will develop over the coming months. RM discussed the issue with Steve Fisher and advised that support of R-GMA is required whilst APEL is dependent on it. RM reported that he has spoken to SF and there is currently no change to the R-GMA situation - process ongoing. 290.19 DB/SP to progress the details of the Project Map over the next few months, cross-checking that all elements are incorporated, including strategic priorities and staffing. To be completed before the next Oversight Committee. The meeting closed at 2:30 pm. The next PMB would be at 1:00 pm on Monday 18 February.

Top of Message | Previous Page | Permalink

JiscMail Tools


RSS Feeds and Sharing


Advanced Options


Archives

November 2017
October 2017
September 2017
August 2017
May 2017
April 2017
March 2017
February 2017
January 2017
October 2016
August 2016
July 2016
June 2016
May 2016
April 2016
March 2016
February 2016
January 2016
December 2015
November 2015
October 2015
September 2015
July 2015
May 2015
April 2015
March 2015
February 2015
January 2015
December 2014
November 2014
October 2014
September 2014
August 2014
July 2014
June 2014
May 2014
April 2014
March 2014
February 2014
January 2014
December 2013
October 2013
August 2013
July 2013
June 2013
May 2013
March 2013
February 2013
January 2013
December 2012
November 2012
October 2012
September 2012
August 2012
July 2012
June 2012
May 2012
April 2012
March 2012
February 2012
January 2012
December 2011
November 2011
October 2011
September 2011
August 2011
July 2011
June 2011
May 2011
April 2011
March 2011
February 2011
January 2011
December 2010
November 2010
October 2010
September 2010
August 2010
July 2010
June 2010
May 2010
April 2010
March 2010
February 2010
January 2010
December 2009
November 2009
October 2009
September 2009
August 2009
July 2009
June 2009
May 2009
April 2009
March 2009
February 2009
January 2009
December 2008
November 2008
October 2008
September 2008
August 2008
July 2008
June 2008
May 2008
April 2008
March 2008
February 2008
January 2008
December 2007
November 2007
October 2007
September 2007
August 2007
July 2007
June 2007
May 2007
April 2007
March 2007
February 2007
January 2007
2006
2005
2004
2003
2002
2001
2000


WWW.JISCMAIL.AC.UK

Secured by F-Secure Anti-Virus CataList Email List Search Powered by the LISTSERV Email List Manager